Mike's salary is $2000 more than fred together they make $5800 find both salaries.
Answer by ikleyn(52798)

.

Mike          + Fred = 5800      (1)

(Fred + 2000) + Fred = 5800


2*Fred = 5800 - 2000


2*Fred = 3800


  Fred = 3900/2 = 1900 dollars (Fred's salary)

  From equation (1),  Mike = 5800 - Fred = 5800 - 1900 = 3900 dollars (Mike's salary).

Solved.

--------------

Is everything clear to you in my solution ?